mj23:

Traci spent $45.20 in 5 hours of shopping at the mall. $ ________ per hour

mj23:

@Ultrilliam

harliii:

you have to divide 45.20/5 and u will get ur answer srry not allow to give out exact answer

mj23:

its ok and thx for the help

harliii:

yw

mj23:

Solve this problem. Find 75% of 375. Type the correct answer rounded to two decimals.

harliii:

100%/x%=375/75 (100/x)*x=(375/75)*x - we multiply both sides of the equation by x 100=5*x - we divide both sides of the equation by (5) to get x 100/5=x

mj23:

thx I got 1 more then I'm done

mj23:

A whole number consists of the digits 2, 4, 6, 8, and 0. If each digit is used only once, how many whole numbers are possible? Type the correct answer. ________

harliii:

5 digits 5*4*3*2*1=120 ways if you count 0 in front of the number as a whole number if you don't, then there are 96 ways

mj23:

ok thanks so much

harliii:

yw

mj23:

I got 100%

harliii:

yw :)